The Morris Corporation has $450,000 of debt outstanding, and it pays an interest rate of 8% annually. Morris's annual sales are $2.25 million, its average tax rate is 40%, and its net profit margin on sales is 3%. If the company does not maintain a TIE ratio of at least 4 to 1, its bank will refuse to renew the loan and bankruptcy will result. What is Morris's TIE ratio? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

TIE ratio is Times Interest Earned

Formula is , TIE ratio=

EBIT=Earnings before interest and taxes

I=interest payments

Let's calculate these values

I=interest payments= Outstanding debt X interest rate=$450,000 X 8%=$36,000

So, I= $36,000

Now, EBIT= Net Profit + Interest Payments (I) + Tax paid

Net Profit= Sales X NetProfit Margin = $2,250,000 X 3% = $67,500

I= $36,000

Tax Paid=

where t = tax rate=40%=0.4

So, Tax Paid= ($67,500 X 0.4) /(1-0.4)= ($67,500 X 0.4)/0.6 = $27,000 / 0.6 = $ 45,000

Now, EBIT = $67,500 + $36,000 + $45,000= $148,500

TIE ratio= $148,500/$36,000= 4.125 = ~4.12 (rounding off to two decimal places)

Since the TIE ratio is better (more) than 4, bank will NOT declare bankruptcy.

Instead, the bank will accept to issue new loan.

Please rate my answer if it helped you. Thank you.